OpenStudy (anonymous):

I will give medal and fan, but I need explanation. You apply a horizontal force of 10.0 N to pull a wheeled suitcase at a constant speed of 0.5m/s across a flat ground. How much power is used? I know that power is equal to work/time, but I don't know how to get the time.

OpenStudy (ipwnbunnies):

You don't need all that fancy stuff. You can use it for the derivation, but it includes calculus I believe. Power = Force*Velocity Voila.

OpenStudy (anonymous):

Yes, Power = Force * Velocity. iPwnBunnies, You don't need calculus if both the force and velocity are constant because they both come outside of the integral. Source: https://en.wikipedia.org/wiki/Power_%28physics%29#Mechanical_power

OpenStudy (shamim):

Power p=w/t=F*s/t=F*v=10*0.5=?

OpenStudy (shamim):

Unit of power is watt=w